# Robust Radiotherapy Planning with Spatially Based Uncertainty Sets

### v1

@@ 5.2 Robust Model Evaluation
Next, the spatially bound uncertainty set $(\mathcal{U}_{\text{SB}})$, introduced in Section 2.2, will be constructed and customized for our patient data. Subsequently, the benefits of both the robust model (2.2.2) based on this uncertainty set, and the dose-volume methodology of Section 4, will be examined and discussed.
@@ 5.2.1 Uncertainty Set Construction
The method of constructing our uncertainty set is justified in Section 2.2 using the data of brain cancer Patient 1 from the TCIA archive in Clark et al. (2013), illustrated in Figure 1. The graphs in Figure 4, which are based on the Brain data of Table 1 (Patient 4 from the archive), demonstrate a similar pattern to the one that emerges from Figure 1. Both figures clearly illustrate that the radiosensitivity distance increases in the geometric voxel distance, up to a geometric distance of approximately ten. To estimate the difference in radiosensitivity between voxel pairs as a function of the voxel distance $\Delta$, we adopt the following model with parameters $\alpha_{0}$, $\alpha_{1}$, $\alpha_{2}$, and $\gamma$ (initially 0),
@@ 5.2.2 Robust Formulation Experiments
We now evaluate optimal solutions of spatially robust model (2.2.2) and compare them with optimal solutions of the nominal formulation (2.1) for the Brain dataset. We also compare with the robust box model that is a robust model (2.2) under a box uncertainty set ($\mathcal{U}_{B}$). The solutions to the robust box model can also be obtained by solving (2.2.2) with $\underaccent{\bar}{\phi}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}=\underaccent{\bar}{\phi}^{0}}$ and $\bar{\phi}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}=\bar{\phi}^{0}}$ by setting $\gamma_{uv}$ in the $\mathcal{U}_{\text{SB}}$ uncertainty set to a large value (it suffices to set $\gamma_{uv}=1$) for all $u,v\in T$. Thus, the robust box case can be viewed as a special (simple) case of our suggested approach.
Optimal solutions of (2.2.2) (with $\mathcal{U}_{\text{SB}}$) are computed for a range of uncertainty set parameter values, $\delta=0.02,0.04,\ldots,0.14$ and $\gamma=0.01,\ldots,0.05$, and a range of homogeneity parameter values, $\mu=1.0625,1.0750,\ldots,1.4000$. Note that for some combinations of uncertainty set and homogeneity parameters, formulation (2.2.2) has $x=0$ as the only feasible solution; such solutions are omitted from the graphs. For the box uncertainty set ($\mathcal{U}_{B}$), the solution were computed with the same values for $\delta$, however, a larger range of homogeneity parameters, $\mu=1.0625,1.0750,\ldots,1.4750$, is explored to be able to obtain nonzero solutions for all $\delta$ values. Moreover, in order to obtain a more accurate value of $\mu$ below which there is no nonzero dose solution, for each $(\gamma,\delta)$ pair, the $\mu$ values between the first nonzero dose solution and the last zero dose solution were refined, first in increments of $0.0025$, then in incremenets of $0.0005$, and finaly in increments of $0.0001$. To evaluate a plan involving beamlet decision variable vector $x^{*}$ under uncertainty set $\mathcal{U}$, which could be a singleton when considering the nominal case, empirical measures of worst-case dose and homogeneity are defined as follows
@@
Before comparing the performance of solutions that are optimal to (2.2.2) under different scenarios, it should be noted that the optimal solution value of (2.2.2) decreases in the uncertainty set parameters $\delta$ and $\gamma$ (see Figure 13 in Appendix D). Note that combinations of $\mu$, $\delta$, and $\gamma$ resulting in a zero beamlet intensity configuration ($x={\mathbf{0}}$) are omitted from all figures and are interpreted as “infeasible” plans.
@@
Figure 5: Nominal biologically-adjusted performance of solutions optimal to model (2.2.2) for $\delta=0.14$ and different values of $\gamma$, compared to the nominal solution.
@@
A “cost-of-robustness” can be measured by evaluating a robust solution in the nominal scenario where there is no uncertainty. Figures 5 and 6 illustrate this phenomenon. The nominal formulation (2.1) corresponds to setting $\delta=0$ and $\gamma$ sufficiently large in $\mathcal{U}_{\text{SB}}$; for the particular dataset being considered (Patient 4), the nominal formulation is equivalent to a setting of $\delta=0$ and $\gamma\geq 0.04$ in formulation (2.2.2)). On the other hand, uncertainty sets with $\gamma\leq 0.03$ exclude the nominal scenario. Consequently, models with smaller $\gamma$ values tend to perform poorly, at least in terms of the minimum biological dose, in the nominal scenario. Figure 5 shows the performance in the nominal scenario, for $\delta=0.14$, with different values of $\gamma$ and $\mu$ in (2.2.2) (a curve is presented for each value of $\gamma$, obtained by varying the values of $\mu$). In general, the “cost of robustness” appears more evident as $\gamma$ decreases: the minimal $\hat{\mu}$ that is attained with a nonzero dose solution increases, and $\hat{\underaccent{\bar}{d}}$ somewhat decreases. This can be explained by the fact that the nominal scenario is excluded from the uncertainty set when $\gamma\leq 0.03$, and the distance between this scenario and the uncertainty set increases as $\gamma$ decreases. The figure also shows that the robust box solution and the robust spatially bound solution with $\gamma=0.04$ exhibit fairly similar performance in the nominal scenario but the robust box appears less stable and demonstrates a notable drop in dose for biological homogeneity values slightly larger than 1.1. Note that a lack of stability stems from the fact that the measured $\hat{\mu}$ and $\hat{\underaccent{\bar}{d}}$ values typically differ from the values of $\mu$ and $\hat{\underaccent{\bar}{d}}$, respectively, that are used in optimizing (2.2.2). Consequently, even slight changes in the parameter $\mu$ may in some cases induce a different optimal solution that is associated with significantly different values of the measured $\hat{\mu}$ and $\hat{\underaccent{\bar}{d}}$. In this case, when increasing $\mu$ from $1.3603$ to $1.3607$, the nominal performance of the robust box model solution improves with respect to both homogeneity and dose, from $\hat{\mu}=1.105$ and $\hat{\underaccent{\bar}{d}}=37.4$ to $\hat{\underaccent{\bar}{\mu}}=1.1$ and $\hat{\underaccent{\bar}{d}}=55.1$.
@@
Figure 6 shows nominal performance curves for $\gamma=0.02,0.04,1$, $\delta=0.02,0.04,\ldots,0.1{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}4}$, and a range of $\mu$ values (each curve is obtained by varying the value of $\mu$). For each of the $\gamma$ values, the “cost of robustness” becomes more evident as $\delta$ increases: the curves for larger $\delta$ values appear to be nearly dominated in the $\hat{\mu}-\hat{\underaccent{\bar}{d}}$ space by the curves obtained for smaller $\delta$ values. The robust box results appear much less stable than the corresponding SR model results for $\gamma\in\{0.02,0.04\}$, especially for values of $\hat{\mu}\leq 1.11$, with doses 25Gy lower than the optimal nominal solution compared to at most 5Gy lower for $\gamma=0.02$ and $0.04$. Recall that solutions of (2.2.2) with $\gamma=0.02$ are not designed to protect against the nominal realization (which is excluded from the uncertainty sets with $\gamma<0.04$). Hence, the solutions obtained in the case where $\gamma=0.02$ have a biologically-adjusted dose of approximately $0.81$ Gy less than the corresponding optimal nominal dose for all $\hat{\mu}>1.11$. When increasing $\gamma$ to $0.04$, the uncertainty set captures the nominal solution and so the dose obtained for $\hat{\mu}\geq 1.11$ become equal to those of the nominal model, albeit the minimal $\hat{\mu}$ for which there was a nonzero dose is higher than the nominal model, even for small values of $\delta$. This trend persists when moving to the robust box model, which results in almost optimal nominal performance for $\delta\leq 0.04$.
@@
Figure 6: Nominal biologically-adjusted performance of solutions optimal to models (2.1) and (2.2.2). (for $\gamma=0.02$, $\delta=0$ is infeasible.)
Next, we consider the physical dose performance of the obtained solutions. Although physical performance measures do not account for the variable radiosensitivity, and planning based on such measures may not utilize the information that has been recently made available by advanced biomarkers such as FMISO, they are still important as the ongoing and current basis for treatment guidelines and clinical practice. Keeping in mind that physical performance is not explicitly maximized in the optimization formulations (2.1) or (2.2.2), it may be interesting to examine how the resulting planning solutions fare with respect to the physical performance measures. Figure 7 compares solutions of (2.2.2) with $\delta=0.14$ and different values of $\gamma$, the robust box solution (also with $\delta=0.14$), and the nominal solution. Figure 6A depicts the physical homogeneity vs. minimal physical dose of the different models, and Figure 6B shows the EUD (equivalent uniform dose). The EUD measure was introduced as a single number representing the dose throughout the PTV (Niemierko 1997, Choi and Deasy 2002), and is given by the formula $EUD=\left(\frac{1}{|T|}\sum_{v\in T}d_{v}^{\alpha}\right)^{\frac{1}{\alpha}},$ where, as in Nohadani and Roy (2017) we use $\alpha=-10$. Surprisingly, both the optimal nominal solution and the robust box solution exhibit poor performance compared with the solutions of SR model (2.2.2) with respect to both physical homogeneity and physical minimal PTV dose, for physical $\mu$ values in the interval $(1.07,1.10)$. Specifically, the solutions of the nominal and robust box models are less stable and, with both physical dose and EUD up to 20Gy lower than for $gamma\leq 0.05$, and a minimal physical homogeneity value of around $1.07$ compares to $1.02$ for $\gamma\leq 0.05$. Similar behavior can be observed for other values of $\delta$ tested; see Figure 15 in Appendix D. A possible explanation for this behavior is that under both the nominal and the box uncertainty plans, the dose is allocated in a way that is focused mainly on the PTV voxels with the lowest radiosensitivities (following the max-min formulation). The spatially bound uncertainty may smooth out the PTV radiosensitivity uncertainty, resulting in a worst case that considers a greater number of voxels, including voxels for which the radiosensitivity is nominally higher, thus increasing the minimal physical dose.
@@ B EUD
Figure 7: Physical performance of solutions optimal to (2.2.2) with $\delta=0.14$ and a range of $\gamma$ values, compared to the nominal solution.
@@ D Box
Figure 8: (7A) – PET, pink and red levels indicating low oxygen. (7B)-(7D) – physical dose.
@@
Figures 8-9 display the physical dose obtained by the nominal solution, the robust box solution with $\delta=0.14$, and the robust spatial solution of (2.2.2) with $(\delta,\gamma)=(0.14,0.04)$. For each model, the value of the homogeneity parameter $\mu$ is chosen so that it yields the highest minimal physical dose while maintaining physical homogeneity less than 1.1 (which is a standard clinical requirement). All three dose images in Figure 7B-7D are aligned with the same PET/CT FMISO scan slice 7A. The robust box and nominal solutions apply wider physical dose ranges to the PTV, 61.5-67.5 Gy for the robust box and 62-68 Gy for the nominal, compared to the robust spatial solution, with a range of 63-65 Gy. Furthermore, the nominal and robust box solutions create dose “hot spots” of approximately 100 Gy in the perimeter of the exterior OAR, while the robust spatial model solution applies a maximal physical dose of less than 71 Gy to this OAR. This may be due to the fact that the nominal and robust box solutions use fewer beams, with some of the beams set to a higher intensity, than the robust spatial solution. Thus, while the robust spatial model retains approximately the same adjusted dose as the nominal solution, it accounts for the variability and uncertainty in the radiosensitivities. It also appears to avoid excessively high OAR doses.
@@
Figure 10: Worst-case biologically-adjusted performance ($\hat{\mu}$, $\hat{\underaccent{\bar}{d}}$) of solutions optimal to model (2.2.2) with misspecified $\gamma$, under the assumption that the “true” $\mathcal{U}_{\text{SB}}$ has $\delta=0.1$ and $\gamma$ as indicated. The performance of the nominal solution is also shown for comparison.
Finally, to evaluate the cost of misspecifying the uncertainty set parameters, Figures 10 and 11 depict the worst-case performance of the various solutions under the assumption that the correct uncertainty models differ with respect to the parameter values assumed in formulation (2.2.2). In Figure 10 the robust solutions use the “true value” of $\delta=0.1$, so the misspecification is only in terms of $\gamma$. Specifically, $\gamma=0.02$ and $\gamma=0.04$ in Figure 9A and Figure 9B, respectively. Note that while the worst-case performance deteriorates as the value of $\gamma$ used to derive the solution deviates from the true value, the dose declines by at most (and is generally much less than) $1.1$ Gy. Importantly, the spatially robust models may achieve homogeneity values that are close to the homogeneity of the correct model, improving the homogeneity by more than $0.05$ relative to the nominal and robust box solutions. This is evident for most values of $\delta$ (see Figures 16 and 17 in Appendix D). Moreover, although the robust box and nominal model appear to attain a similar biologically-adjusted dose as the true model for high values of $\hat{\mu}$, the performance of these models appears much less stable for desirable (lower) values of $\hat{\mu}$. Notably, the robust box obtains doses significantly lower than all models considered including the nominal one for some values of $\hat{\mu}$. Hence, even when $\gamma$ is misspecified in (2.2.2), it may still yield nearly the same homogeneity, $\hat{\mu}$, perhaps at a slight cost to $\hat{\underaccent{\bar}{d}}$.
@@
Figure 11: Worst-case biologically-adjusted performance ($\hat{\mu}$, $\hat{\underaccent{\bar}{d}}$) of solutions optimal to (2.2.2) with misspecified $\delta$, under the assumption that the “true” $\mathcal{U}_{\text{SB}}$ has $\gamma=0.04$ and $\delta$ as indicated
Figure 11 depicts experimental results for evaluating the effect of misspecifying $\delta$ in the uncertainty model. In this set of experiments, the true uncertainty set, as well as the uncertainty sets assumed in solving (2.2.2), all share the same value of $\gamma=0.04$, but differ in the values of $\delta$. While the misspecification of $\delta$ seems to cause only a modest decrease in the value of $\hat{\underaccent{\bar}{d}}$ (of up to $0.5$ Gy) compared with the value obtained under the true $\delta$, while the homogeneity, $\hat{\mu}$, appears to be more significantly affected. Specifically, underestimating $\delta$ increases $\hat{\mu}$ of up to $0.03$ (for a similar dose), while overestimating $\delta$ may lead to little or no improvement in the homogeneity (a small decrease in $\hat{\mu}$). Further, the $\hat{\mu}$ values attained for the nominal model are significantly worse (larger than the spatially robust models by up to 0.06) for the higher true value of $\delta=0.14$, compared to being approximately 0.03 larger than all spatially robust models when the true $\delta=0.02$.

### v2

@@ 5.2 Robust Model Evaluation
Next, the spatially bound uncertainty set $(\mathcal{U}_{\text{SB}})$, introduced in Section 2.2, will be constructed and customized for our patient data. Subsequently, the benefits of both the robust model (10) based on this uncertainty set, and the dose-volume methodology of Section 4, will be examined and discussed.
@@ 5.2.1 Uncertainty Set Construction
The method of constructing our uncertainty set is justified in Section 2.2 using the data of brain cancer Patient 1 from the TCIA archive in Clark et al., (2013), illustrated in Figure 1. The graphs in Figure 4, which are based on the Brain data of Table 1 (Patient 4 from the archive), demonstrate a similar pattern to the one that emerges from Figure 1. Both figures clearly illustrate that the radiosensitivity distance increases in the geometric voxel distance, up to a geometric distance of approximately ten. To estimate the difference in radiosensitivity between voxel pairs as a function of the voxel distance $\Delta$, we adopt the following model with parameters $\alpha_{0}$, $\alpha_{1}$, $\alpha_{2}$, and $\gamma$ (initially 0),
@@ 5.2.2 Robust Formulation Experiments
We now evaluate optimal solutions of spatially robust model (10) and compare them with optimal solutions of the nominal formulation (1) for the Brain dataset. We also compare with the robust box model that is a robust model (3) under a box uncertainty set ($\mathcal{U}_{B}$). The solutions to the robust box model can also be obtained by solving (10) with $\underaccent{\bar}{\phi}{=\underaccent{\bar}{\phi}^{0}}$ and $\bar{\phi}{=\bar{\phi}^{0}}$ by setting $\gamma_{uv}$ in the $\mathcal{U}_{\text{SB}}$ uncertainty set to a large value (it suffices to set $\gamma_{uv}=1$) for all $u,v\in T$. Thus, the robust box case can be viewed as a special (simple) case of our suggested approach.
Optimal solutions of (10) (with $\mathcal{U}_{\text{SB}}$) are computed for a range of uncertainty set parameter values, $\delta=0.02,0.04,\ldots,0.14$ and $\gamma=0.01,\ldots,0.05$, and a range of homogeneity parameter values, $\mu=1.0625,1.0750,\ldots,1.4000$. Note that for some combinations of uncertainty set and homogeneity parameters, formulation (10) has $x=0$ as the only feasible solution; such solutions are omitted from the graphs. For the box uncertainty set ($\mathcal{U}_{B}$), the solution were computed with the same values for $\delta$, however, a larger range of homogeneity parameters, $\mu=1.0625,1.0750,\ldots,1.4750$, is explored to be able to obtain nonzero solutions for all $\delta$ values. Moreover, in order to obtain a more accurate value of $\mu$ below which there is no nonzero dose solution, for each $(\gamma,\delta)$ pair, the $\mu$ values between the first nonzero dose solution and the last zero dose solution were refined, first in increments of $0.0025$, then in incremenets of $0.0005$, and finaly in increments of $0.0001$. To evaluate a plan involving beamlet decision variable vector $x^{*}$ under uncertainty set $\mathcal{U}$, which could be a singleton when considering the nominal case, empirical measures of worst-case dose and homogeneity are defined as follows
@@
Before comparing the performance of solutions that are optimal to (10) under different scenarios, it should be noted that the optimal solution value of (10) decreases in the uncertainty set parameters $\delta$ and $\gamma$ (see Figure 13 in Appendix D). Note that combinations of $\mu$, $\delta$, and $\gamma$ resulting in a zero beamlet intensity configuration ($x={\mathbf{0}}$) are omitted from all figures and are interpreted as “infeasible” plans.
@@
Figure 5: Nominal biologically-adjusted performance of solutions optimal to model (10) for $\delta=0.14$ and different values of $\gamma$, compared to the nominal solution.
@@
A “cost-of-robustness” can be measured by evaluating a robust solution in the nominal scenario where there is no uncertainty. Figures 5 and 6 illustrate this phenomenon. The nominal formulation (1) corresponds to setting $\delta=0$ and $\gamma$ sufficiently large in $\mathcal{U}_{\text{SB}}$; for the particular dataset being considered (Patient 4), the nominal formulation is equivalent to a setting of $\delta=0$ and $\gamma\geq 0.04$ in formulation (10)). On the other hand, uncertainty sets with $\gamma\leq 0.03$ exclude the nominal scenario. Consequently, models with smaller $\gamma$ values tend to perform poorly, at least in terms of the minimum biological dose, in the nominal scenario. Figure 5 shows the performance in the nominal scenario, for $\delta=0.14$, with different values of $\gamma$ and $\mu$ in (10) (a curve is presented for each value of $\gamma$, obtained by varying the values of $\mu$). In general, the “cost of robustness” appears more evident as $\gamma$ decreases: the minimal $\hat{\mu}$ that is attained with a nonzero dose solution increases, and $\hat{\underaccent{\bar}{d}}$ somewhat decreases. This can be explained by the fact that the nominal scenario is excluded from the uncertainty set when $\gamma\leq 0.03$, and the distance between this scenario and the uncertainty set increases as $\gamma$ decreases. The figure also shows that the robust box solution and the robust spatially bound solution with $\gamma=0.04$ exhibit fairly similar performance in the nominal scenario but the robust box appears less stable and demonstrates a notable drop in dose for biological homogeneity values slightly larger than 1.1. Note that a lack of stability stems from the fact that the measured $\hat{\mu}$ and $\hat{\underaccent{\bar}{d}}$ values typically differ from the values of $\mu$ and $\hat{\underaccent{\bar}{d}}$, respectively, that are used in optimizing (10). Consequently, even slight changes in the parameter $\mu$ may in some cases induce a different optimal solution that is associated with significantly different values of the measured $\hat{\mu}$ and $\hat{\underaccent{\bar}{d}}$. In this case, when increasing $\mu$ from $1.3603$ to $1.3607$, the nominal performance of the robust box model solution improves with respect to both homogeneity and dose, from $\hat{\mu}=1.105$ and $\hat{\underaccent{\bar}{d}}=37.4$ to $\hat{\underaccent{\bar}{\mu}}=1.1$ and $\hat{\underaccent{\bar}{d}}=55.1$.
@@
Figure 6 shows nominal performance curves for $\gamma=0.02,0.04,1$, $\delta=0.02,0.04,\ldots,0.1{4}$, and a range of $\mu$ values (each curve is obtained by varying the value of $\mu$). For each of the $\gamma$ values, the “cost of robustness” becomes more evident as $\delta$ increases: the curves for larger $\delta$ values appear to be nearly dominated in the $\hat{\mu}-\hat{\underaccent{\bar}{d}}$ space by the curves obtained for smaller $\delta$ values. The robust box results appear much less stable than the corresponding SR model results for $\gamma\in\{0.02,0.04\}$, especially for values of $\hat{\mu}\leq 1.11$, with doses 25Gy lower than the optimal nominal solution compared to at most 5Gy lower for $\gamma=0.02$ and $0.04$. Recall that solutions of (10) with $\gamma=0.02$ are not designed to protect against the nominal realization (which is excluded from the uncertainty sets with $\gamma<0.04$). Hence, the solutions obtained in the case where $\gamma=0.02$ have a biologically-adjusted dose of approximately $0.81$ Gy less than the corresponding optimal nominal dose for all $\hat{\mu}>1.11$. When increasing $\gamma$ to $0.04$, the uncertainty set captures the nominal solution and so the dose obtained for $\hat{\mu}\geq 1.11$ become equal to those of the nominal model, albeit the minimal $\hat{\mu}$ for which there was a nonzero dose is higher than the nominal model, even for small values of $\delta$. This trend persists when moving to the robust box model, which results in almost optimal nominal performance for $\delta\leq 0.04$.
@@
Figure 6: Nominal biologically-adjusted performance of solutions optimal to models (1) and (10). (for $\gamma=0.02$, $\delta=0$ is infeasible.)
Next, we consider the physical dose performance of the obtained solutions. Although physical performance measures do not account for the variable radiosensitivity, and planning based on such measures may not utilize the information that has been recently made available by advanced biomarkers such as FMISO, they are still important as the ongoing and current basis for treatment guidelines and clinical practice. Keeping in mind that physical performance is not explicitly maximized in the optimization formulations (1) or (10), it may be interesting to examine how the resulting planning solutions fare with respect to the physical performance measures. Figure 7 compares solutions of (10) with $\delta=0.14$ and different values of $\gamma$, the robust box solution (also with $\delta=0.14$), and the nominal solution. Figure 7A depicts the physical homogeneity vs. minimal physical dose of the different models, and Figure 7B shows the EUD (equivalent uniform dose). The EUD measure was introduced as a single number representing the dose throughout the PTV (Niemierko, 1997 and Choi and Deasy, 2002), and is given by the formula $EUD=\left(\frac{1}{|T|}\sum_{v\in T}d_{v}^{\alpha}\right)^{\frac{1}{\alpha}},$ where, as in Nohadani and Roy, (2017) we use $\alpha=-10$. Surprisingly, both the optimal nominal solution and the robust box solution exhibit poor performance compared with the solutions of SR model (10) with respect to both physical homogeneity and physical minimal PTV dose, for physical $\mu$ values in the interval $(1.07,1.10)$. rSpecifically, the solutions of the nominal and robust box models are less stable and, with both physical dose and EUD up to 20Gy lower than for $gamma\leq 0.05$, and a minimal physical homogeneity value of around $1.07$ compares to $1.02$ for $\gamma\leq 0.05$. Similar behavior can be observed for other values of $\delta$ tested; see Figure 15 in Appendix D. A possible explanation for this behavior is that under both the nominal and the box uncertainty plans, the dose is allocated in a way that is focused mainly on the PTV voxels with the lowest radiosensitivities (following the max-min formulation). The spatially bound uncertainty may smooth out the PTV radiosensitivity uncertainty, resulting in a worst case that considers a greater number of voxels, including voxels for which the radiosensitivity is nominally higher, thus increasing the minimal physical dose.
@@ B EUD
Figure 7: Physical performance of solutions optimal to (10) with $\delta=0.14$ and a range of $\gamma$ values, compared to the nominal solution.
@@ D Box
Figure 8: (8A) – PET, pink and red levels indicating low oxygen. (8B)-(8D) – physical dose.
@@
Figures 8-9 display the physical dose obtained by the nominal solution, the robust box solution with $\delta=0.14$, and the robust spatial solution of (10) with $(\delta,\gamma)=(0.14,0.04)$. For each model, the value of the homogeneity parameter $\mu$ is chosen so that it yields the highest minimal physical dose while maintaining physical homogeneity less than 1.1 (which is a standard clinical requirement). All three dose images in Figure 8B-8D are aligned with the same PET/CT FMISO scan slice 8A. The robust box and nominal solutions apply wider physical dose ranges to the PTV, 61.5-67.5 Gy for the robust box and 62-68 Gy for the nominal, compared to the robust spatial solution, with a range of 63-65 Gy. Furthermore, the nominal and robust box solutions create dose “hot spots” of approximately 100 Gy in the perimeter of the exterior OAR, while the robust spatial model solution applies a maximal physical dose of less than 71 Gy to this OAR. This may be due to the fact that the nominal and robust box solutions use fewer beams, with some of the beams set to a higher intensity, than the robust spatial solution. Thus, while the robust spatial model retains approximately the same adjusted dose as the nominal solution, it accounts for the variability and uncertainty in the radiosensitivities. It also appears to avoid excessively high OAR doses.
@@
Figure 10: Worst-case biologically-adjusted performance ($\hat{\mu}$, $\hat{\underaccent{\bar}{d}}$) of solutions optimal to model (10) with misspecified $\gamma$, under the assumption that the “true” $\mathcal{U}_{\text{SB}}$ has $\delta=0.1$ and $\gamma$ as indicated. The performance of the nominal solution is also shown for comparison.
Finally, to evaluate the cost of misspecifying the uncertainty set parameters, Figures 10 and 11 depict the worst-case performance of the various solutions under the assumption that the correct uncertainty models differ with respect to the parameter values assumed in formulation (10). In Figure 10 the robust solutions use the “true value” of $\delta=0.1$, so the misspecification is only in terms of $\gamma$. Specifically, $\gamma=0.02$ and $\gamma=0.04$ in Figure 10A and Figure 10B, respectively. Note that while the worst-case performance deteriorates as the value of $\gamma$ used to derive the solution deviates from the true value, the dose declines by at most (and is generally much less than) $1.1$ Gy. Importantly, the spatially robust models may achieve homogeneity values that are close to the homogeneity of the correct model, improving the homogeneity by more than $0.05$ relative to the nominal and robust box solutions. This is evident for most values of $\delta$ (see Figures 16 and 17 in Appendix D). Moreover, although the robust box and nominal model appear to attain a similar biologically-adjusted dose as the true model for high values of $\hat{\mu}$, the performance of these models appears much less stable for desirable (lower) values of $\hat{\mu}$. Notably, the robust box obtains doses significantly lower than all models considered including the nominal one for some values of $\hat{\mu}$. Hence, even when $\gamma$ is misspecified in (10), it may still yield nearly the same homogeneity, $\hat{\mu}$, perhaps at a slight cost to $\hat{\underaccent{\bar}{d}}$.
@@
Figure 11: Worst-case biologically-adjusted performance ($\hat{\mu}$, $\hat{\underaccent{\bar}{d}}$) of solutions optimal to (10) with misspecified $\delta$, under the assumption that the “true” $\mathcal{U}_{\text{SB}}$ has $\gamma=0.04$ and $\delta$ as indicated
Figure 11 depicts experimental results for evaluating the effect of misspecifying $\delta$ in the uncertainty model. In this set of experiments, the true uncertainty set, as well as the uncertainty sets assumed in solving (10), all share the same value of $\gamma=0.04$, but differ in the values of $\delta$. While the misspecification of $\delta$ seems to cause only a modest decrease in the value of $\hat{\underaccent{\bar}{d}}$ (of up to $0.5$ Gy) compared with the value obtained under the true $\delta$, while the homogeneity, $\hat{\mu}$, appears to be more significantly affected. Specifically, underestimating $\delta$ increases $\hat{\mu}$ of up to $0.03$ (for a similar dose), while overestimating $\delta$ may lead to little or no improvement in the homogeneity (a small decrease in $\hat{\mu}$). Further, the $\hat{\mu}$ values attained for the nominal model are significantly worse (larger than the spatially robust models by up to 0.06) for the higher true value of $\delta=0.14$, compared to being approximately 0.03 larger than all spatially robust models when the true $\delta=0.02$.